# Semantic Web technologies in sensor-based personal health monitoring systems: A systematic mapping study

### v2

@@ Abstract
In recent years, there has been an increased focus on early detection, prevention, and prediction of diseases. This, together with advances in sensor technology and the Internet of Things, has led to accelerated efforts in the development of personal health monitoring systems. This study analyses the state of the art in the use of Semantic Web technologies in sensor-based personal health monitoring systems. Using a systematic approach, a total of 43 systems are selected as representative of the current state of the art. We critically analyse the extent to which the selected systems address seven key challenges: interoperability, context awareness, situation detection, situation prediction, decision support, explainability, and uncertainty handling. We discuss the role and limitations of Semantic Web technologies in managing each challenge. We then conduct a quality assessment of the selected systems based on the data and devices used, system and components development, rigour of evaluation, and accessibility of research outputs. Finally, we propose a reference architecture to provide guidance for the design and development of new systems. This study provides a comprehensive mapping of the field, identifies inadequacies in the state of the art, and provides recommendations for future research.
Semantic Web technologies,
ontologies,
knowledge graphs,
linked data,
sensors,
Internet of Things,
@@ 1 Introduction
Non-communicable diseases are on the rise globally, resulting not only in decreased quality of life but also increasing healthcare costs [1]. For this reason, there have been accelerated efforts to develop personal health monitoring systems for early detection, prediction, and prevention of diseases. The emerging paradigm of precision health goes beyond treating existing diseases and rather focuses on preventing disease before it strikes. Eschewing the one-size-fits-all approach in favour of assessing individual circumstances, precision health encourages people to actively monitor and work towards improving their health so as to lower the risk of disease [2]. Personal health monitoring is part of this vision, allowing people to not only increase understanding of their health but also to receive recommendations for any necessary interventions. Significant advances in the Internet of Things (IoT) over the last decade has led to the rapid rise of wearable sensors, which are increasingly being used for health monitoring outside traditional clinical settings. Wearable sensors can collect and measure physiological data such as vital signs, which can be combined with health records and questionnaires to determine lifestyle habits and medical history. Additionally, ambient sensors can monitor environmental factors such as air quality and weather, which have a significant impact on health.
@@
There are several crucial issues affecting sensor-based personal health monitoring systems, which can be distilled into seven key challenges. The first of these is interoperability. Heterogeneous sensor observations, differing data transmission technologies, and disparate standards for describing health data all contribute to interoperability issues in personal health monitoring systems. Additionally, the representation of health domain knowledge and its integration with sensor data remains a challenging task [3]. By its nature, sensor data is dynamic and complex, necessitating interpretation into higher-level concepts or situations [4]. Situation analysis involves the use of sensor data to detect the current state of a given environment (situation detection), while anticipating possible future states (situation prediction) [5]. The representation of domain knowledge is essential for facilitating situation analysis from sensor data and supporting subsequent decision-making. The decision support process augments human judgement, assisting clinicians in navigating complex medical decisions [6] and supporting patients in making informed health decisions outside clinical settings [7]. Both the situation analysis and decision support processes must incorporate context awareness. Dey and Abowd [8] define context as any information that can be used to characterize the situation of an entity, including location, identity, activity, and time. Such information is essential for accurate situation analysis and targeted decision support. Moreover, since the health domain is a high-stakes one, explainability is gaining traction as a pivotal aspect of AI-driven health systems [9]. Finally, given the probabilistic nature of health outcomes and the limitations of sensor and other health data, there is inherent uncertainty in the situation analysis and decision-making processes [10]. Thus, effective uncertainty handling is critical in sensor-based personal health monitoring systems.
Semantic Web technologies, which are widely used in the health domain, can alleviate some of these key challenges. The goal of this study is to systematically map the state of the art in the use of Semantic Web technologies in sensor-based personal health monitoring systems. The contributions of this paper are as follows:
@@
    We present a systematic mapping of the field based on 43 systems that are systematically selected as representative of the current state of the art.
@@
    We critically evaluate the extent to which the systems address the seven key challenges, i.e. interoperability, context awareness, situation detection, situation prediction, decision support, explainability, and uncertainty handling. We discuss the role and limitations of Semantic Web technologies in managing each challenge.
@@ 2.1 Sensor-based personal health monitoring
Health monitoring sensors are generally either wearable or implantable. Wearable sensors are worn on the body or are otherwise integrated with clothes and shoes. Such sensors include electrodes for measuring electrical signals, thermal sensors for measuring temperature, and PPG sensors. Smart watches and bands are the most commonly used wearable sensors, but earables (devices placed in the ear) have recently emerged as a promising alternative [15]. In contrast, implantable sensors operate from within the human body. Although they are much less commonly used than wearable sensors, they are particularly useful for monitoring chronic illness as well as post-surgery monitoring to minimise complications and avoid readmission [16]. Health monitoring sensors also include portable devices that can measure physiological and activity data but cannot be practically worn or used for prolonged periods of time. Examples of these include blood pressure monitors and pulse oximeters, as well as smartphones which contain sensors such accelerometers, which measure acceleration, and gyroscopes, which measure orientation and angular velocity [17]. Additionally, ambient sensors are increasingly being incorporated in health monitoring to monitor the state of the external environment, such as temperature, humidity, and air quality, as these factors have a significant impact on human health [12, 18].
@@ 2.2 Semantic Web technologies
Three overlapping Semantic Web technologies have emerged as the most prominent over the years: ontologies, knowledge graphs, and linked data [22].
@@ 2.2.1 Ontologies
As domain-agnostic ontologies, both SSN and SAREF require augmentation to meet application-specific requirements [26]. SAREF provides a suite of ontologies that extend the core ontology for different domains, including two that are relevant for personal health monitoring: SAREF4EHAW [^2] https://saref.etsi.org/saref4ehaw/ for eHealth and ageing well and SAREF4WEAR [^3] https://saref.etsi.org/saref4wear/ for wearable devices. SAREF4EHAW provides support for modelling concepts such as health system actors (including patients and caregivers) and health devices (including wearables), with the wearable concept linked to the SAREF4WEAR ontology. An additional extension, SAREF4Health, was developed to address the limitations of SSN and SAREF in representing real-time ECG time series data exchanged between mobile devices and cloud gateways [25]. In contrast to SSN, SAREF is targeted at industry developers rather than ontology experts [25], making it practical for real-world applications. Furthermore, its extensions for the health domain provide a solid foundation for building semantic personal health monitoring systems. Additional representational support can be obtained by integrating resources such as standardized clinical terminologies and medical knowledge bases.
@@ 2.2.2 Knowledge graphs
A knowledge graph can generally be understood as a knowledge base of real-world data represented in a graph-based data model. Ontologies are a vital building block of knowledge graphs, used to define the data schema (such as properties, restrictions, and relationships) and enable semantic reasoning and entailment [27]. Knowledge graphs have seen increasingly widespread use in the health domain. Their graph structure enables the conceptualisation, representation, and integration of data [27]. This is advantageous in health monitoring systems, where the integration of various sources of health data is critical. An example of this is the Precision Medicine Knowledge Graph (PrimeKG), which integrates diverse biomedical data from multiple sources with the goal of enabling precision medicine analyses [28]. Previous research has also explored the automatic construction of knowledge graphs from electronic health records [29, 30], which can then be used for clinical decision support. Additionally, knowledge graphs have been proposed for health risk prediction [31], drug discovery [32], and as a tool for explainability in AI-driven health systems [33, 34]. Knowledge graphs have also proven useful in sensor-based systems, for example by providing graph-based visualisations of the data generated by IoT devices, which can then be queried in real time [35].
@@ 2.2.3 Linked data
2.2.4 Languages and standards
The development of Semantic Web technologies is facilitated using different languages and standards. Resource Description Framework (RDF) [^4] https://www.w3.org/RDF, a standard for the description and exchange of interconnected data in the form of subject-predicate-object triples, can be considered one of the core building blocks of the Semantic Web. Several extensions to RDF have been proposed. These include RDF Schema (RDFS) [^5] https://www.w3.org/wiki/RDFS, which provides a vocabulary to enrich RDF data; RDF-star [^6] https://w3c.github.io/rdf-star/cg-spec, which allows the subject or object of a triple to refer to another triple; and the Notation 3777https://w3c.github.io/N3/spec specification, which extends the representational abilities of RDF by supportive declarative programming and allowing the access of online knowledge. Other important standards in the Semantic Web community are: eXtensible Markup Language (XML), a markup language and file format; Web Ontology Language (OWL) [^8] https://www.w3.org/OWL, a language for constructing ontologies; Semantic Web Rule Language (SWRL) [^9] https://www.w3.org/Submission/SWRL, a language for expressing rules and logic; Shapes Constraint Language (SHACL) [^10] https://www.w3.org/TR/shacl, a language for describing RDF graphs, which also includes a rules language; and SPARQL Protocol and RDF Query Language (SPARQL) [^11] https://www.w3.org/TR/rdf-sparql-query, a language for retrieving and manipulating RDF data. SPARQL-star extends SPARQL to allow querying and updating of RDF-star data.
@@ 3 Related reviews
Several reviews related to sensors, Semantic Web technologies, and the health domain have been published. These reviews can generally be categorised into three overlapping groups, which are illustrated as a Venn diagram in Figure 1. The reviews in Group 1 focus on the use of Semantic Web technologies in the health domain; those in Group 2 review the use of sensors and IoT in the health domain; those in Group 3 review the use of Semantic Web technologies with sensor and IoT data; and finally, Group 4 consists of other related reviews that do not fit neatly into any of the first three groups. The related reviews are discussed in detail in the remainder of this section and summarised in Table 1.
@@ 3.1 Group 1: Semantic Web technologies in the health domain
This group of reviews explores the use of Semantic Web technologies in healthcare. Zenuni et al. [40] review ontologies and semantic data repositories used in different aspects of the health domain, including hospital systems and health datasets. A similar review is conducted by Haque et al. [9], who explore themes such as e-healthcare, disease diagnosis, and information management. Peng et al. [41] and Hammad et al. [42] focus on semantic approaches for health data integration and management, including data from wearable devices. Dimitrieski et al. [43] review ontologies and ontology alignment approaches in healthcare, while Jing et al. [44] focus on ontologies for rule management in clinical decision support systems. Although the reviews in this group provide a good overview of the ways in which Semantic Web technologies have been used in the health domain, half of them do not mention sensors or IoT at all, while the other half do not include sensor data as a major focus.
@@ 3.2 Group 2: Sensors and IoT in the health domain
This group considers the use of sensors and IoT in the health domain. Islam et al. [45] and Yin et al. [46] conduct general surveys on IoT for healthcare, covering a broad range of considerations on the topic including networks, communication standards and protocols, and cybersecurity. The review by Qi et al. [47] focuses on the use of IoT in personalised healthcare systems, including sensor devices and data processing techniques. Philip et al. [48] explore advances in the field such as cloud computing, while Albahri et al. [49] focus on health monitoring systems for telemedicine applications, highlighting techniques that support the connection of hospital services to remote patients. There have also been reviews specifically focusing on the state of the art in wearable sensors for health monitoring, such as those by Cusack et al. [18], Dias and Cunha [12], and Majumder et al. [50]. Kim et al. [51] hone in on biosensors that detect biofluids, such as sweat and tears, while Baig et al. [52] highlight the potential of remote monitoring systems for clinical adoption. Punj and Kumar [53], Banaee et al. [54], and Andreu-Perez et al. [16] explore advances in wearable sensor data collection, mining, and processing, and Dang et al. [55] focus on statistical analysis and machine learning (ML) as modelling tools. While these reviews provide useful analyses on the role of sensors and IoT in health monitoring, they either do not mention Semantic Web technologies or do so briefly without an in-depth analysis of their role in health monitoring.

### v3

@@ Abstract
In recent years, there has been an increased focus on early detection, prevention, and prediction of diseases. This, together with advances in sensor technology and the Internet of Things, has led to accelerated efforts in the development of personal health monitoring systems. This study analyses the state of the art in the use of Semantic Web technologies in sensor-based personal health monitoring systems. Using a systematic approach, a total of 48 systems are selected as representative of the current state of the art. We critically analyse the extent to which the selected systems address seven key challenges: interoperability, situation detection, situation prediction, decision support, context awareness, explainability, and uncertainty handling. We discuss the role and limitations of Semantic Web technologies in managing each challenge. We then conduct a quality assessment of the selected systems based on the data and devices used, system and components development, rigour of evaluation, and accessibility of research outputs. Finally, we propose a reference architecture to provide guidance for the design and development of new systems. This study provides a comprehensive mapping of the field, identifies inadequacies in the state of the art, and provides recommendations for future research.
@@ 1 Introduction
Non-communicable diseases are on the rise globally, resulting not only in decreased quality of life but also increasing healthcare costs [1]. For this reason, there have been accelerated efforts to develop personal health monitoring systems for early detection, prediction, and prevention of diseases. The emerging paradigm of precision health goes beyond treating existing diseases and rather focuses on preventing disease before it strikes. Eschewing the one-size-fits-all approach in favour of assessing individual circumstances, precision health encourages people to actively monitor and work towards improving their health so as to lower the risk of disease [2]. Personal health monitoring is part of this vision, allowing people to not only increase understanding of their health but also to receive recommendations for any necessary interventions. Significant advances in the Internet of Things (IoT) over the last decade has led to the rapid rise of wearable sensors, which are increasingly being used for health monitoring outside traditional clinical settings. Wearable sensors can collect and measure physiological data such as vital signs, which can be combined with health records and questionnaires to determine lifestyle habits and medical history. Beyond physiological data, ambient sensors can monitor environmental factors such as air quality and weather, which have a significant impact on health. Additionally, the widespread adoption of artificial intelligence (AI) in the health domain has led to personal health monitoring systems becoming increasingly AI-driven. Such systems use techniques such as knowledge representation and reasoning and machine learning to analyse health data and provide actionable insights.
@@
There are several crucial issues affecting sensor-based personal health monitoring systems, which can be distilled into seven key challenges. The first of these is interoperability. Heterogeneous sensor observations, differing data transmission technologies, and disparate standards for describing health data all contribute to interoperability issues in personal health monitoring systems. Additionally, the representation of health domain knowledge and its integration with sensor data remains a challenging task [3]. By its nature, sensor data is dynamic and complex, necessitating interpretation into higher-level concepts or situations [4]. Situation analysis involves the use of sensor data to detect the current state of a given environment (situation detection), while anticipating possible future states (situation prediction) [5]. The representation of domain knowledge is essential for facilitating situation analysis from sensor data and supporting subsequent decision-making. The decision support process augments human judgement, assisting clinicians in navigating complex medical decisions [6] and supporting patients in making informed health decisions outside clinical settings [7]. Both the situation analysis and decision support processes must incorporate context awareness. Dey and Abowd [8] define context as any information that can be used to characterize the situation of an entity, including location, identity, activity, and time. Such information is essential for accurate situation analysis and targeted decision support. Moreover, since the health domain is a high-stakes one, explainability is gaining traction as a pivotal aspect of AI-driven health monitoring systems [9]. Finally, given the probabilistic nature of health outcomes and the limitations of sensor and other health data, there is inherent uncertainty in the situation analysis and decision-making processes [10]. Thus, effective uncertainty handling is critical in sensor-based personal health monitoring systems.
Semantic Web technologies have been used in numerous health applications ranging from health data integration [11, 12] to clinical decision support [13, 14]. While they have shown promise in alleviating the seven key challenges, the degree to which they address each challenge differs. Therefore, the goal of this study is to systematically map the state of the art in the use of Semantic Web technologies in sensor-based personal health monitoring systems. We analyse the effectiveness of the systems in addressing the seven key challenges, identify the role and limitations of Semantic Web technologies, and assess the overall quality of the systems. Accordingly, a systematic mapping study was selected as the most appropriate approach. Systematic mapping studies have become increasingly popular in software engineering [15]. While systematic reviews aim at synthesizing evidence for specific research questions, mapping studies go further and provide a high-level view of the research landscape [15]. By structuring a research area through classification and categorisation, mapping studies seek to discover emerging research trends and identify potential gaps for further lines of inquiry [16, 17].
The contributions of this study are as follows:
@@
    We present a systematic mapping of the field based on 48 systems that are systematically selected as representative of the current state of the art. Our selection criteria includes systems across all development stages, from research prototypes to production-level systems.
@@
    We critically evaluate the extent to which the systems address the seven key challenges, i.e. interoperability, situation detection, situation prediction, decision support, context awareness, explainability, and uncertainty handling.
@@
    We assess the role and limitations of Semantic Web technologies in managing each challenge.
  4. 4.
@@ 2.1 Sensor-based personal health monitoring
Health monitoring sensors are usually either wearable or implantable. Wearable sensors are worn on the body or are otherwise integrated with clothes and shoes. Such sensors include electrodes for measuring electrical signals, thermal sensors for measuring temperature, and PPG sensors. Smart watches and bands are the most commonly used wearable sensors, but earables (devices placed in the ear) have recently emerged as a promising alternative [22]. In contrast, implantable sensors operate from within the human body. Although they are much less commonly used than wearable sensors, they are particularly useful for monitoring chronic illness as well as post-surgery monitoring to minimise complications and avoid readmission [23]. Health monitoring sensors also include portable devices that can measure physiological and activity data but cannot be practically worn or used for prolonged periods of time. Examples of these include blood pressure monitors and pulse oximeters, as well as smartphones which contain sensors such accelerometers, which measure acceleration, and gyroscopes, which measure orientation and angular velocity [24].
Besides wearable and implantable sensors, there exist contact-free sensors that can monitor health-related factors. For example, the commercially available Emfit QS device [^1] https://emfit.com/heart-rate-variability-hrv-during-sleep sensor uses ballistocardiography (BGC), a measure of ballistic forces generated by the heart, to measure heart rate variability during sleep when placed underneath one’s mattress. Additionally, device-free human sensing is also gaining traction in many domains, including health [25]. Examples include the use of radio frequency signal reflections and WiFi channel state information to passively monitor vital signs [26] and detect falls [27] without requiring physical devices. Additionally, ambient sensors are increasingly being incorporated in health monitoring to monitor the state of the external environment, such as temperature, humidity, and air quality, as these factors have a significant impact on human health [19, 28].
@@ 2.2 Semantic Web technologies
The Semantic Web provides foundational mechanisms for knowledge representation and reasoning, thereby playing an important role in the design of AI-driven health monitoring systems. Three overlapping Semantic Web technologies have emerged as the most prominent over the years: ontologies, knowledge graphs, and linked data [32]. We will begin with an overview of key Semantic Web languages and standards, after which we will discuss each of the three technologies in turn, highlighting how they contribute to health monitoring.
2.2.1 Languages and standards
The development of Semantic Web technologies is facilitated using different languages and standards. Resource Description Framework (RDF) [^2] https://www.w3.org/RDF, a standard for the description and exchange of interconnected data in the form of subject-predicate-object triples, can be considered one of the core building blocks of the Semantic Web. As RDF is an abstract data model, it can be serialised in different formats, including N-Triples, Terse RDF Triple Language (Turtle), eXtensible Markup Language (XML), and JavaScript Object Notation for Linked Data (JSON-LD) [33]. Several extensions to RDF have been proposed. These include RDF Schema (RDFS) [^3] https://www.w3.org/wiki/RDFS, which provides a vocabulary to enrich RDF data; RDF-star [^4] https://w3c.github.io/rdf-star/cg-spec, which allows an RDF triple to be embedded as the subject or object of another triple, without necessarily asserting the embedded triple, thereby enabling richer metadata annotation; and the Notation3555https://w3c.github.io/N3/spec specification, which extends the representational abilities of RDF by supportive declarative programming and allowing the access of online knowledge. Other important standards in the Semantic Web community are: Web Ontology Language (OWL) [^6] https://www.w3.org/OWL, a language for constructing ontologies; Semantic Web Rule Language (SWRL) [^7] https://www.w3.org/Submission/SWRL, a language for expressing rules and logic; Shapes Constraint Language (SHACL) [^8] https://www.w3.org/TR/shacl, a language for describing RDF graphs, which also includes a rules language; and SPARQL Protocol and RDF Query Language (SPARQL) [^9] https://www.w3.org/TR/rdf-sparql-query, a language for retrieving and manipulating RDF data. SPARQL-star extends SPARQL to allow querying and updating of RDF-star data, while SPARQL Inferencing Notation (SPIN) [^10] https://spinrdf.org is a rules language based on SPARQL.
@@ 2.2.2 Ontologies
As domain-agnostic ontologies, both SSN and SAREF require augmentation to meet application-specific requirements [37]. SAREF provides a suite of ontologies that extend the core ontology for different domains, including two that are relevant for personal health monitoring: SAREF4EHAW [^12] https://saref.etsi.org/saref4ehaw for eHealth and ageing well and SAREF4WEAR [^13] https://saref.etsi.org/saref4wear for wearable devices. SAREF4EHAW provides support for modelling concepts such as health system actors (including patients and caregivers) and health devices (including wearables), with the wearable concept linked to the SAREF4WEAR ontology. An additional extension, SAREF4Health, was developed to address the limitations of SSN and SAREF in representing real-time ECG time series data exchanged between mobile devices and cloud gateways [36]. In contrast to SSN, SAREF is targeted at industry developers rather than ontology experts [36], making it more readily adoptable by those without extensive ontology development experience. Furthermore, its extensions for the health domain provide a solid foundation for building semantic personal health monitoring systems. Additional representational support can be obtained by integrating resources such as standardized clinical terminologies and medical knowledge bases.
@@ 2.2.3 Knowledge graphs
A knowledge graph can generally be understood as a knowledge base of real-world data represented in a graph-based data model. Ontologies are a vital building block in many knowledge graphs, and are used to define their data schema (such as properties, restrictions, and relationships) as well as enable semantic reasoning and entailment [38]. Knowledge graphs have seen increasingly widespread use in the health domain. Their graph structure enables the conceptualisation, representation, and integration of data [38]. This is advantageous in health monitoring systems, where the integration of various sources of health data is critical. An example of this is the Precision Medicine Knowledge Graph (PrimeKG), which integrates diverse biomedical data from multiple sources with the goal of enabling precision medicine analyses [39]. Previous research has also explored the automatic construction of knowledge graphs from electronic health records [40, 41], which can then be used for clinical decision support. This is related to personal health knowledge graphs, which are used to represent and reason over individual health data, including data from sensors [42] and electronic health records [43, 44]. Additionally, knowledge graphs have been proposed for drug discovery [45] and as a tool for explainability in AI-driven health monitoring systems [46, 47]. Knowledge graphs have also proven useful in sensor-based systems, for example by providing graph-based visualisations of the data generated by IoT devices, which can then be queried in real time [48].
@@ 3 Related reviews
Several reviews related to sensors, Semantic Web technologies, and the health domain have been published. These reviews can generally be categorised into three overlapping groups, which are illustrated as a Venn diagram in Figure 1. The reviews in Group 1 focus on the use of Semantic Web technologies in the health domain; those in Group 2 review the use of sensors and IoT in the health domain; those in Group 3 review the use of Semantic Web technologies with sensor and IoT data; and finally, Group 4 consists of other related reviews that do not fit neatly into any of the first three groups. The related reviews are discussed in detail in the remainder of this section and summarised in Table LABEL:related-reviews-table.
@@ 3.1 Group 1: Semantic Web technologies in the health domain
This group of reviews explores the use of Semantic Web technologies in healthcare. Zenuni et al. [53] review ontologies and semantic data repositories used in different aspects of the health domain, including hospital systems and health datasets. A similar review is conducted by Haque et al. [9], who explore themes such as e-healthcare, disease diagnosis, and information management. Peng et al. [11] and Hammad et al. [12] focus on semantic approaches for health data integration and management, including data from wearable devices. Dimitrieski et al. [54] review ontologies and ontology alignment approaches in healthcare, while Jing et al. [13] focus on ontologies for rule management in clinical decision support systems. More recently, Amar et al. [55] examine semantic interoperability issues in electronic health records based on the Fast Healthcare Interoperability Resources (FHIR) standard, highlighting how RDF and OWL can improve interoperability. The review by Miranda et al. [56] explores how Semantic Web technologies can be used to enhance the interoperability and management of electronic health records in healthcare systems. Similarly, Wu et al. [57] analyse recent work at the intersection of Semantic Web technologies and electronic health records, with a particular focus on how such technologies can improve data quality. Finally, Cui et al. [14] provide a comprehensive review of healthcare knowledge graphs, discussing their construction and use in a wide range of health applications, including clinical decision support and pharmaceutical research. Although the reviews in this group provide a good overview of the ways in which Semantic Web technologies have been used in the health domain, five of them do not mention sensors at all, while the remaining five reviews do not include sensor data as a major focus.
@@ 3.2 Group 2: Sensors and IoT in the health domain
This group considers the use of sensors and IoT in the health domain. Islam et al. [58] and Yin et al. [59] conduct general surveys on IoT for healthcare, covering a broad range of considerations on the topic including networks, communication standards and protocols, and cybersecurity. The review by Qi et al. [60] focuses on the use of IoT in personalised healthcare systems, including sensor devices and data processing techniques. Philip et al. [61] explore advances in the field such as cloud computing, while Albahri et al. [62] focus on health monitoring systems for telemedicine applications, highlighting techniques that support the connection of hospital services to remote patients. There have also been reviews specifically focusing on the state of the art in wearable sensors for health monitoring, such as those by Babu et al. [63], Cusack et al. [28], Dias and Cunha [19], and Majumder et al. [64]. Kim et al. [65] hone in on biosensors that detect biofluids, such as sweat and tears, while Baig et al. [66] highlight the potential of remote monitoring systems for clinical adoption. Punj and Kumar [67], Banaee et al. [68], and Andreu-Perez et al. [23] explore advances in wearable sensor data collection, mining, and processing, and Dang et al. [69] focus on statistical analysis and machine learning (ML) as modelling tools. Bollineni et al. [70] adopt a forward-looking perspective in their review, highlighting both emerging technologies and future prospects for health-IoT architectures. While these reviews provide useful analyses on the role of sensors and IoT in health monitoring, ten of them do not mention Semantic Web technologies, while the remaining six do so briefly without an in-depth analysis of their role in health monitoring.